40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5860     Entry time: Thu Nov 10 05:54:23 2011
Author: kiwamu 
Type: Update 
Category: Green Locking 
Subject: Beat-note detected : PSL vs Y arm 

[Katrin / Kiwamu]
The beat-note between the PSL green laser and the Y end green laser was successfully detected.
The detection was done by the new broad-band RFPD.
The next step will be an extraction of the frequency fluctuation signal using the delay-line-mixer frequency discriminator.

 

Here is a picture of the RF spectrum analyzer displaying the direct output signal from the broad-band RFPD.
The beat-note was moving around 100 MHz with an RF power of -36 dBm. The frequency fluctuation was about +/- 7MHz in a time scale of 1 sec or so.
DSC_3606_small.jpg


(What we did)
 + Connected a BNC cable which goes from the c1iscey's DAC to the laser slow input
    => this enables a remote control of the laser frequency via the temeperature actuation
 + Realigned the beam pointing of the Y end green laser
 + Installed all the necessary optics on the PSL table
     => currently the PSL green light is adjusted to completely S-polarization
 + readjusted the mode matching telescopes
     => the Y green beam becomes the one with a long Rayleigh range
 + Health check on the broad-band RFPD to see if it is working
 + Installed the BB-RFPD with a +/-15V power supply
 + Fine alignment of the beam combining path
 + Fine tuning of the Y end laser temperature
     => T_PSL = 31.72 deg when the slow FSS feedback is zero.
     => Based on Bryan's measurement (see #elog) the Y end laser temperature was adjusted to 34.0 deg by applying an offset to the slow input.
 + Found the beat note at 100 MHz or so.
     => optimizing the alignment of the beam combining path by maximizing the peak height of the beat-note.
     => maximum peak height observed with an RF spectrum analyzer was about -36 dBm.

ELOG V3.1.3-